MATH 131 – Calculus For Physical and Life Sciences, section 1 Problem Set Guidelines Fall 2007

Weekly homework assignments will be posted each Friday on the course homepage and collected the following Friday. Here are some guidelines to follow while doing your assignments, for our mutual benefit:

• Please be neat! Write up the problems in order. Use only one side of the page (don’t worry too much about saving paper), and staple your assignment together.

• Work on making your solutions clear and concise. Make good use of notation and diagrams.

• Leave space for our grader to write comments.

• Show enough detail so that a confused student in the class could follow your solution. Don’t expect that the reader already knows how to solve the problem. Just writing a number or formula is almost never sufficient.

• When a problem asks for an explanation, you should provide one written in complete sentences. Proofread what you have said to be sure it makes sense.

• No “and then a miracle occurs” moments in solutions, please! If you are left with a gap in a calculation that you just can’t bridge, acknowledge it. The grader will indicate how to fill the gap, if appropriate, in comments.

• Start the assignment early (e.g. the day it is posted). This will allow you ample time to consult with me or the Calculus Workshop, or a study group if you get stuck on some problems. If you start the assignment on the day before it is due, you will be very unhappy!

• If you work with other students to solve a problem, you must write up your final solutions independently. (Plan on taking scrap notes when you work out the problem first, and then “writing it up” in a clean form by yourself afterwards.) Add a note to your solution listing the other people you consulted.